aakkk:

At a summer camp, one large tub of mayonnaise is used for each lunch service. The function below represents the amount of mayonnaise, in tablespoons, left in the tub after N students are served lunch. M(N) = 4000 - 4N What is the domain of M(N)?

Extrinix:

So to solve for \(\sf{N}\), we can remove \(\sf{M(N)}\) and make it into \(\sf{0}\). \(\sf{0=4000-4N}\) Simply solve for \(\sf{N}\). \(\sf{0=4000-4N}\) Subtract 4000 from each side. \(\sf{-4000=-4N}\) Divide both sides by -4. \(\sf{\dfrac{-4000}{-4}=N}\) Simplify it. \(\sf{1000=N}\) And you have what N is! Now, that given, you have to find the domain. This can be done by simply finding 2 points along the line that can represent the graph (in this case, we start with 4000 tablespoons at 0 students and go down by 4 per each student. Let's use the easiest one we can do, (1000,0), the amount of mayonnaise left in the jars. Given these two points: (0,4000) , (1000,0) we can find both interval notation AND inequality notation. Interval notation: \(\sf{[0, 1000]}\) - Both sides are closed (hence the brackets []), which mean that the numbers 0 and 1000 are included Inequality notation: \(\sf{0 \le x \le 1000}\) - Both sides are closed (hence the \(\sf{\le}\) symbol, which means the same as interval notation. "Zero is less than or equal to x which is less than or equal to 1000." That's as in-depth as I can currently explain it considering it's 03:00. Best of luck, \(\style{font Family:Arial; text-shadow: 2px 2px 3px white, 0 0 10px white, 0 0 5px white;color:#23B7F3;font-size:100px;padding:50px;width:500px;height:100%;position:absolute;left:0;z-index:500;text-align:center;} {\Huge\sf\color {} {{-~Akuma~matata}}}\) "what a wonderful phrase"